Use the method of substitution to solve the system of linear equations. 3a=7-b and 3a= 5-b. There is one solution. What is the solution?
Answer by stanbon(75887)

Use the method of substitution to solve the system of linear equations. 3a=7-b and 3a= 5-b. There is one solution. What is the solution?
------
3a = 7-b
3a = 5-b
-----
Substitute for "3a" to get:
7-b = 5-b
Note: That is a contradiction.
There is no solution.
----------
Cheers,
Stan H.
------------